### v2

# João Pereira1,2, Vasco Lopes2,3, David Semedo1, João Neves3

Abstract
Large Vision-Language Models (LVLMs) demonstrate remarkable performance in short-video tasks such as video question answering, but struggle in long-video understanding. The linear frame sampling strategy, conventionally used by LVLMs, fails to account for the non-linear distribution of key events in video data, often introducing redundant or irrelevant information in longer contexts while risking the omission of critical events in shorter ones. To address this, we propose Self-ReS, a non-linear spatiotemporal self-reflective sampling method that dynamically selects key video fragments based on user prompts. Unlike prior approaches, Self-ReS leverages the inherently sparse attention maps of LVLMs to define reflection tokens, enabling relevance-aware token selection without requiring additional training or external modules. Experiments demonstrate that Self-ReS can be seamlessly integrated into strong base LVLMs, improving long-video task accuracy and achieving up to 46% faster inference speed within the same GPU memory budget. The code can be found here: https://github.com/jaca-pereira/Self-ReS.
Index Terms:
Large Vision-Language Models, Video Understanding
I Introduction
Recently, Large Vision-Language Models (LVLMs) emerged as a way to leverage the strong zero- and few-shot performance of Large Language Models (LLMs) in downstream vision tasks. These models combine vision and language by projecting visual features from a vision encoder into the LLM’s embedding space, followed by fine-tuning on multimodal data. In the video domain, the adoption of strong backbone models [1][2] and the development of high-quality pre-training and instruction-tuning datasets [3] has led to improved performance on tasks such as video question answering [4]. While LVLMs demonstrate strong capabilities in understanding short videos, their performance on longer videos remains a significant challenge, due to difficulties in modeling larger numbers of frames. In the reference benchmark Video-MME [4], several state-of-the-art models such as LLaVA-Video [3] report strong performance for short videos, but struggle to obtain competitive performance for longer inputs, especially without the textual hints present in subtitles.
Efforts to improve performance for longer videos can be categorized into two main approaches: a) increasing context size, and b) applying token compression techniques. Often, models propose a combination of the two. Despite significantly increasing the number of sampled frames - e.g., up to 768 for Qwen2-VL [5] and 1024 for NVILA [6]- and applying aggressive token compression to minimize the cost of reasoning over larger contexts, these models still struggle with long videos, being unable to outperform models with significantly smaller context sizes - e.g., LLaVA-Video with 64 frames only
Figure 1: Linear Frame Sampling vs Self-Reflective Sampling (Self-ReS). The linear frame sampling method, used by most LVLMs, samples a dense or uniform subset of frames, increasing the likelihood of sampling redundant and background information or missing key events. In contrast, Self-ReS increases sampling granularity, optimizing context-window usage with prompt-relevant fragments of the video identified through spatiotemporal self-reflection.
We argue that current frame sampling strategies overlook video non-linearity, where critical events, such as a sudden action or a key object appearance, occur in a concentrated spatiotemporal region, rather than being evenly distributed throughout the video - e.g., Figure 1. Moreover, linearly scaling the number of frames to accommodate longer videos will likely introduce background and redundant information, inefficiently allocating more computational resources.
An important line of research focuses on proposing improved sampling strategies for video LVLMs. Widely adopted techniques such as token pooling [7] [6] or merging [8] allow for extended context sizes, but also fail to account for the non-linear nature of video data. More recent approaches attempt to address this non-linearity through dynamic memory banks [9], adaptive token compression  [10][11] or by learning a summarization token [12]. Notably, the proposed solutions rely on external, trainable modules, introducing added computational requirements that may affect their adoption in real-world scenarios.
In this paper, we introduce Self-ReS, a novel Self-Reflective Sampling approach designed to enhance LVLM performance on long video understanding. Self-ReS addresses video non-linearity by selecting key spatiotemporal fragments, guided by textual cues present in the user prompt. Our approach builds on the idea that LLMs can self-reflect and critique their own generations [13] to improve their final answer. In particular, Self-ReS leverages the inherent self-reflective ability of base LLMs to identify relevant video fragments according to a user-specified task.
Self-ReS introduces independent reflection paths, where each path processes, in parallel, a different segment of the video. These paths converge at specific depths within the LLM, enabling unified inference across the entire video. The self-reflective nature of the LLM is expressed in reflection tokens, which serve as judges to evaluate the relevancy of visual tokens based on the user query. By leveraging the inherent ability of LVLMs to generate sparse attention maps[14], which naturally focus on a subset of visual tokens, Self-ReS defines reflection tokens to prioritize critical information. Once the reflection paths converge, the LLM performs inference using its default context size, concentrating on the most salient visual tokens, which are non-linearly selected according to the textual cues in the user prompt. With this strategy, Self-ReS achieves up to $\mathbf{46}\textbf{\%}$ faster inference speed and higher accuracy than the base model while maintaining GPU memory budget.
Our contributions can be summarized as follows:
  • We propose Self-ReS, a Self-Reflective Sampling method that dynamically identifies and prioritizes key spatiotemporal fragments of video data. By leveraging the inherent self-reflective capabilities of LLMs, Self-ReS addresses spatiotemporal non-linearity without requiring additional training or external modules.
  • Self-ReS enables LVLMs to process significantly more frames while reducing computational overhead. It achieves higher accuracy with up to 46% faster inference speed and identical GPU memory usage, making it practical for resource-constrained scenarios.
  • Using Self-ReS, LLaVA-Video outperforms existing open and closed-source models on two reference video benchmarks.
II Related Work
II-A Large Vision-Language Models
LVLMs extend the zero-shot and instruction-following capabilities of LLMs to visual inputs, enabling multimodal tasks across images and videos. Foundational LVLMs such as LLaVA [15] and BLIP [16] connect vision encoders to LLMs via linear projections or learned queries, forming the basis for pioneer video-specific models such as VideoChatGPT [7]. These models are pre-trained on video-captioning and instruction-tuning datasets [3], enabling strong performance in short video tasks.
Recent advancements scale LVLMs to longer videos using state-of-the-art backbones [1][2] and increased frame context sizes [6][5]. However, these improvements often come at the cost of significant computational overhead. Moreover, these methods fail to address spatiotemporal non-linearity in video data, where key events are concentrated in small regions rather than distributed uniformly. Self-ReS enables foundational models to process longer videos within the default context size, dynamically prioritizing salient tokens while maintaining computational efficiency.
II-B Video Sampling Strategies for LVLMs
Sampling strategies are critical for extending LVLM performance to longer videos. Early approaches such as VideoChatGPT [7] and ChatUniVi [8] employed token pooling or merging to handle the larger number of video frames. Memory-based strategies such as MovieChat [9] store long context video information in memory tokens. LLaMA-ViD [17] learns a single token per frame. Recent proposals include learnable video summarization tokens in VideoXL [12] and spatiotemporal adaptive token compression in LongVU [10], which uses high-resolution vision encoders to compute frame similarity and cross-modal relevance. While effective to an extent, these methods often rely on additional backbones and require extensive training [10][12][17], increasing computational costs. Moreover, they may fail to adequately address video non-linearity [9][7][8], where key events occur in sparse spatiotemporal fragments of the video. In contrast, Self-ReS leverages the LLM’s self-reflective abilities, expressed in the form of native attention weights that dynamically identify and prioritize critical visual tokens. Our plug-and-play approach eliminates the need for external modules or additional training, offering a computationally efficient solution that is faster than the baseline and surpasses prior methods in accuracy.
III Methodology
In this section, we detail the methodology behind Self-ReS, illustrated in Figure 2. Self-ReS is designed to address three key challenges faced by LVLMs in long video understanding: 1) Non-Linearity of video data, where critical events, such as short actions or key object appearances, are often concentrated in small spatial and temporal regions of a video, rather than being evenly distributed; 2) Limited context sizes which restrict the number of frames that can be processed, risking the omission of key events in shorter temporal windows; 3) Linearly increasing context size introduces background noise and redundant information, leading to inefficient resource utilization.
Self-ReS addresses these challenges through a process called spatiotemporal self-reflection. This process begins by sampling a large number of frames from a long video and distributing them across parallel reflection paths. Each path processes a segment of the video independently up to a specified layer depth in the LLM. The paths are then merged into a single inference path, where the most salient visual tokens are non-linearly selected based on the user task. By dynamically prioritizing key information and minimizing redundancy, Self-ReS effectively expands the context size without inefficiently scaling computational resources.
III-A Spatiotemporal Self-Reflection
Inspired by the concept that LLMs can critique their own predictions through parallel and independent inferences [13], we design reflection tokens to judge the relevancy of visual tokens for a given task. Our approach leverages the observation that the last input token of LVLMs inherently generates sparse attention maps for visual tokens [14]. By adopting the last input token as the reflection token, Self-ReS avoids the need for additional training.
Without loss of generality, we assume a standard LVLM architecture composed of a vision encoder $f_{V}$, a linear projection matrix $W\in\mathbb{R}^{d_{V}\times d}$, and an LLM $f$, where $d_{V}$ and $d$ correspond to the visual encoder and LLM token dimensionality, respectively. We also assume the default input sequence formatting, with a system prompt $X_{\text{sys}}$ being concatenated with the video features, followed by the user question $X_{q}$.
To address the context size limitations, Self-ReS emulates a larger context size by sampling $T$ frames from a long video and distributing them across $R$ independent reflection paths. Each path processes $S=\frac{T}{R}$ frames, where $S$ matches the base model’s default context size. The frames are first processed by the Vision Encoder $f_{V}$, generating visual features $X_{V}\in\mathbb{R}^{R\times N_{V}\times d_{V}}$, where $N_{V}$ represents the number of visual tokens per reflection path. These features are then projected into the LLM embedding space using the matrix $W$, producing $\hat{X}_{V}\in\mathbb{R}^{R\times N_{V}\times d}$. The inputs for each reflection path $r_{i}$ are defined as:
$X_{r_{i}}=[X_{\text{sys}}\oplus\hat{X}_{V}^{i}\oplus X_{q}],$ (1)
with $X_{r_{i}}\in\mathbb{R}^{N_{r}\times d}$, where $\hat{X}_{V}^{i}$ corresponds to the projected visual tokens of path $r_{i}$, $N_{r}$ is the total number of tokens per reflection path, and $\oplus$ denotes concatenation.
Each reflection path sequence $X_{r_{i}}$ is forwarded independently, thus having its own attention mask and position embeddings. Depending on the available memory resources, Self-ReS can operate in parallel (batch mode) or sequentially (memory-efficient mode), ensuring scalability across diverse computational environments. For simplicity, we assume parallel inference mode in this explanation.
Our self-reflective approach begins with the LLM’s prefill stage, performing a forward pass up to a designated self-reflection layer. At this layer, Self-ReS applies Self-Reflective Attention (SRA), a modified version of Scaled Dot-Product Attention designed to perform a non-linear selection of visual tokens. This mechanism determines query-key similarities without the memory-intensive value multiplication and softmax operations. Also, introducing a new attention mechanism only at the self-reflection layer allows compatibility with all existing LVLM setups, including those using Flash Attention [18], which typically does not expose attention weights.
Specifically, we designate the hidden embedding of the last token of the user query $X_{q}$, at self-reflection layer $l$, as our reflection token. The reflection token is then projected as the query vector $q_{l}\in\mathbb{R}^{N_{V}\times d_{q}}$. Likewise, the hidden embeddings of the visual tokens are projected into the key matrix $K_{V_{l}}\in\mathbb{R}^{N_{V}\times d_{k}}$, where $d_{k}=d_{q}$. SRA is computed as:
$\text{SRA}(q_{l},K_{V_{l}})=\frac{q_{l}\cdot K^{T}_{V_{l}}}{\sqrt{d_{k}}}.$ (2)
We purposely omit the attention head dimension for simplicity. The final saliency scores are computed by averaging the scores for each token over all attention heads, isolating individual contributions.
Figure 2: Illustration of Self-ReS methodology. Self-ReS splits the long-context video into reflection paths, performing batch inference to a specific depth. Reflection tokens then assess the relevancy of visual tokens for the user task. After self-reflection, the paths merge into a single inference path using either Regular or Smooth Inference modes. Regular mode samples salient tokens and restarts the forward pass from the first layer, while Smooth mode continues from the self-reflection layer with the most salient hidden embeddings.
The last stage of spatiotemporal self-reflection utilizes the judgments produced by SRA to evaluate the relevancy of visual tokens across reflection paths. These judgments are derived from the scores calculated from the sparse attention maps of the reflection tokens. The expressiveness of these scores arises from the fact that reflection tokens $q_{l}$ act as highly informational embeddings for next-token prediction [14].
Once the attention weights are calculated, Self-ReS selects the top-$N_{V}$ visual tokens:
$S_{c}=\text{Top}_{N_{v}}\left(\{\text{SRA}_{r_{i}}(q_{l},K_{V_{l}})\}_{r_{i}=1}^{r_{i}=R}\right),$ (3)
where $S_{c}$ contains the indices of the most salient visual tokens over all reflection paths. This selection process introduces non-linearity in both the spatial and temporal dimensions and eliminates background information, retaining only the most critical tokens. The stored indices in $S_{c}$ are then used to sample tokens in the next phase, where the reflection paths converge into a unified inference path.
III-B Converging Reflection Paths
The final stage of Self-ReS merges the reflection paths into a single inference path, enabling the model to operate within its default context size while reasoning over the non-linear salient video fragments selected during spatiotemporal self-reflection. This step allows the model to effectively capture relationships across video segments. To achieve this, Self-ReS implements two complementary strategies for converging reflection paths, tailored to balance computational efficiency and full model reasoning.
Regular Inference
This mode initiates a new forward pass from the first layer using only the top-$N_{v}$ visual tokens selected during the spatiotemporal self-reflection phase. The input sequence is constructed as:
$X=[X_{\text{sys}}\oplus\hat{X}_{r}\oplus X_{\text{q}}],$ (4)
where $\hat{X}_{r}\in\mathbb{R}^{N_{v}\times d}$ represents the set of projected visual tokens with the highest scores $S_{c}$, from the flattened sequence of visual tokens across all reflection paths. The original temporal order of the visual tokens is maintained, enabling the LLM to perform a full inference pass on the selected salient tokens.
Smooth Inference
This mode leverages the hidden embeddings after the Self-Reflection layer $l$, and continues the forward pass with the top-$N_{v}$ visual tokens, instead of restarting from the first layer. We define the input sequence for the final inference path as:
$H=[H_{\text{sys}_{0}}\oplus H_{r}\oplus H_{\text{q}_{0}}],$ (5)
where $H_{\text{sys}_{0}}$ and $H_{\text{q}_{0}}$ are the hidden embeddings of the system and user tokens from the first reflection path, and $H_{r}\in\mathbb{R}^{N_{v}\times d}$ represents the top-$N_{v}$ self-reflected hidden embeddings of salient visual tokens, conditioned on the user query across all reflection paths. As shown in Figure 3, the hidden embeddings of text tokens generated with the same prompt across different reflection paths converge to equivalent representations. We wield this insight to define our smooth converging strategy and select the textual tokens of a single reflection path. Smooth Inference minimizes redundant computations, maintaining high inference efficiency.
Final Inference Path
Regardless of the chosen strategy, Self-ReS converges into a unified new inference path sequence. This sequence is forwarded using vanilla attention and reassigned positional embeddings (see section -A of the Appendix). This path focuses exclusively on the most salient visual tokens, identified through the LLM’s self-reflective capabilities guided by the user query. By leveraging these tokens, the final model prediction effectively captures global inter-segment dependencies, enabling reasoning over a larger context of the input video. Notably, this is achieved without actually increasing the model’s default context size or memory requirements, ensuring efficient handling of non-linear video events.
III-C Complexity of Self-ReS
Our proposed SRA effectively computes saliency scores for visual tokens, similar to regular attention mechanisms but with reduced computational overhead. In our baseline LVLM, the computational cost of attention is quadratic over the total number of input tokens - $O(R*N_{r}^{2})$, where $R$ is the number of reflection paths and $N_{r}$ is the number of input tokens per reflection path. In Self-ReS, while attention computation remains quadratic, it is computed over a smaller number of input tokens split between reflection paths: $O(R*N_{r}^{2})=O(R*(N/R)^{2})=O(N^{2}/R)$. Thus, while Self-ReS maintains quadratic complexity, it is reduced by a factor of $R$. SRA (Eq. 2) has linear complexity $O(N_{v})$, which is then applied to each reflection path $O(R\cdot N_{V})$. Therefore, it does not affect the overall complexity of the LVLM inference flow.
Figure 3: Consistency of system and textual hidden representations over reflection paths. The hidden representations of the system and textual input have been reduced to a low-dimensional space (2D), evidencing the residual difference between the hidden representations over different reflection paths.
IV Experiments
In this section, we thoroughly evaluate the impact of integrating Self-ReS in a state-of-the-art foundational LVLM using two reference video understanding benchmarks. We benchmark Self-ReS against existing foundational LVLMs and state-of-the-art sampling methods to demonstrate its effectiveness in improving accuracy and efficiency.
IV-A Datasets.
Video-MME [4]
Encompasses multiple choice questions over challenging short, medium and long videos. The videos cover a wide range of topics including Knowledge, Film & Television, Sports Competition, Artistic Performance, Life Record, and Multilingual. We focus on the benchmark without subtitles because it closely simulates real-world scenarios where subtitles are unavailable, and ensures that the evaluation better reflects the model’s ability to comprehend visual spatiotemporal information.
MLVU [19]
The Multi-task Long Video Understanding benchmark is used to evaluate performance of the proposed strategy on diverse and complex long-video understanding tasks. MLVU comprises a collection of videos ranging from 3 minutes to 2 hours, spanning diverse genres such as movies, documentaries, and surveillance footage. The benchmark features nine distinct task categories, including topic reasoning, anomaly recognition, summarization, and action understanding, designed to test holistic, single-detail, and multi-detail video comprehension.
IV-B Implementation Details
For all experiments, we use a greedy decoding strategy. These values ensure deterministic and reproducible results. For Self-ReS specific parameters, we explore different combinations of total frames sampled ($T$), segment size ($S$), and the self-reflection layer depth, as detailed in the ablation study. We rely on an open-source library [20] to implement SRA at a designated self-reflection layer and batch processing for reflection paths.
IV-C Long Video Understanding
Table I compares the performance of Self-ReS against other LVLMs with different sampling strategies. Results clearly demonstrate the superiority of Self-ReS in comparison to methods using similar sized LLMs, achieving state-of-the-art results. LLaVA-Video equipped with Self-ReS surpasses both the base model and models with much larger context sizes. Additionally, our spatiotemporal self-reflective sampling methodology helps LLaVA-Video outperform much larger closed and open source methods, including GPT-4o [21] and GPT-4o mini [22].
These results evidence the benefit of non-linear token selection at both the spatial and temporal dimensions of the video. Self-ReS uses non-linear spatiotemporal self-reflection to compress task-relevant information into a significantly smaller context window, contrasting with current approaches that linearly extend the context across numerous frames, leading to redundancy and background noise. Furthermore, our training-free approach, leveraging the naturally sparse attention maps of the last user query token, allows for our method to surpass other trainable, non-linear sampling methods, by building on and improving a strong base model.
TABLE I: Performance comparison on Video-MME and MLVU.
Model Sampling $\bm{T}$ Video-MME MLVU
Strategy Long All
Large >=>=> =72B VideoLLaMA2[23] Q-Former 32 57.6 62.4 45.6
LLaVA-OV[24] Pooling 32 60.0 66.3 68.0
GPT-4o mini[22] - 250 58.6 64.8 -
GPT-4o[21] - 384 65.3 71.9 64.6
Mid-Sized 7B MovieChat[9] Memory 2048 - - 18.0
LLaMA-ViD[17] CrossAt 1fps - - 17.2
ChatUniVI[8] Merging 64 35.8 40.6 -
VideoXL[12] VST 2048 55.5 61.0 64.9
LongVU[10] SVA 1fps - - 65.4
Oryx[11] CrossAt 128 - 58.8 67.5
Qwen2-VL[5] MLP 738 - 63.3 -
NVILA[6] Pooling 1024 54.8 64.2 70.1
LLaVA-Video[3] Pooling 32 52.2 62.5 66.5
64 53.4 64.3 70.3
128 54.1 64.6 70.3
LLaVA-Video Self-ReS 128 (32) 54.6 64.3 67.9
128 (64) 56.3 65.5 70.7
$\bm{T}$ denotes the total number of frames sampled for input. VST is Video Summarization Token, SVA is Spatiotemporal Visual Aggregation and CrossAt is Cross-Attention. ’-’ means value not reported.
IV-D Ablation Study
This section analyzes the impact of Self-ReS’s parameters on accuracy. Figure 4 reveals the effect of the depth of the self-reflection layer in the performance of Self-ReS over long videos, using the Video-MME (Long) [4] benchmark. For consistency, we fix the segment size $S$ to $32$ and the total number of sampled frames $T$ to $64$.
We focus our ablation study in the shallow layers of the LLM. Considering the role of different layers in the forward pass of an LLM, early layers typically capture syntactic or input-level features, which are better for predictive tasks [25] such as our designated spatiotemporal self-reflection, aiming to predict scores for visual tokens. We leave most deeper layers of the LLM out of our study, as they provide semantic and abstract representations geared towards text generation.
Our results demonstrate that the Regular Inference strategy, where the model resets the forward pass after self-reflection, observes a steady behavior throughout different depths. However, our analysis reports peak accuracy for Smooth Inference at depth 5, and a rapid decline in deeper layers, corroborating the claims in [25]. Since the forward pass continues after self-reflection, and the visual hidden embeddings from different reflection paths are concatenated, the hidden embeddings of earlier layers represent more input-specific representations, while later layers represent abstract features that are relative for the answer being generated for each reflection path, and therefore are both worse for predicting visual token scores and for path convergence. Regular Inference is more robust to this factor, since the forward pass restarts from the first layer using input tokens.
We also conduct additional experiments in Video-MME (Long) using varied configurations of $S$ and $T$, with a fixed self-reflection layer depth of $5$ for optimal performance-efficiency trade-off. The results are provided in Table 4 and indicate a steady gain in accuracy by emulating a larger context size of $T$ frames and consequently more reflection paths $R$. This experiment confirms that we can effectively emulate larger context sizes within the base model’s default context window by performing non-linear sampling.
$S$ $T$ $R$ Video-MME
Long All
32 64 2 53.3 63.2
96 3 53.9 63.6
128 4 54.6 64.3
64 128 2 56.3 65.5
Figure 4: Ablation study on self-reflection layer depth (left) and context size width (right). Experiments reveal that optimal self-reflection performance is achieved in earlier layers. Emulating larger context sizes with increased $T$ values also benefits accuracy.
IV-E Computational Efficiency Analysis
This section evaluates the impact of Self-ReS on the inference speed of the baseline LLaVA-Video [3] model. We measure these metrics on 70 videos from the Video-MME [4] benchmark, after a warm-up round of 30 videos, using a setup of 16GB of RAM and one NVIDIA A100 GPU with 40GB of VRAM.
Figure 5 plots the accuracy in Video-MME (Long) and frames per second (fps) for Self-ReS and the base model across different context sizes. For Self-ReS, we fix the segment size $S$ to 32 and 64, and vary the number of reflection paths $R$, emulating a larger context size of $T$ frames. We use Self-ReS in batch mode, ensuring a fair comparison to the baseline. We purposely omit GPU memory consumption since we observed similar values when using batch mode. For information on inference speed and GPU memory consumption when using the memory efficient sequential mode, we refer the reader to the supplementary material.
We observe up to $\mathbf{46}$faster inference speed and $\mathbf{2.2}$higher accuracy compared to the base model when using an emulated context size of $T=128$. As a result of our non-linear spatiotemporal sampling method, our solution can compress task-relevant information into a much smaller context size. Across all setups, Self-ReS achieves similar or superior accuracy in faster inference speed.
IV-F Qualitative Evaluation
Our proposed spatiotemporal self-reflection process, guided by the user prompt, can effectively focus on the small spatiotemporal window where the main event occurs. In Figure 1, we see that Self-ReS allows the model to correctly infer that the man blowtorches the door handles, despite this event happening in a small spatial and temporal window of the video. Other sampling strategies that perform linear sampling can introduce noise - e.g., frames that are not related to the man or the door handles - or miss key details - e.g., frames where the man interacts with the door lock and not the handles.
Figure 5: Comparison of accuracy and inference speed on Video-MME for base model with and without Self-ReS, across varying context sizes. Results show evidence that Self-ReS has a positive effect on both the efficiency and accuracy of the base model.
V Conclusion
In this paper, we introduce Self-ReS, a novel self-reflective sampling method tailored to address the challenges of long-video understanding in LVLMs. By leveraging the inherent self-reflective abilities of LLMs, Self-ReS selects key spatiotemporal fragments guided by an user query, effectively addressing video non-linearity without requiring additional training or external modules. Experiments on Video-MME and MLVU benchmarks demonstrate that Self-ReS achieves state-of-the-art results among mid-sized LVLMs, surpassing models with significantly larger context sizes while reducing base model inference speed by up to 46%. Self-ReS can be seamlessly integrated into existing LVLM architectures, enabling more efficient and accurate video understanding. In future work, we plan to assess whether Self-ReS can cope with augmented textual contexts and elicit chain-of-thought reasoning.